OpenStudy (waheguru):

ratio question if there were 393000 live births in canada and the atio of boys to girls was 21:20 how may boys and girls were born plz show me how to get the answer DONT TELL ME THE ANSWER BUT HOW TO GET THE ANSWER THANKYOU

OpenStudy (anonymous):

21:20 lets suppose tht 21 boys and 20 girls were born total so total is 41 so solve it by ration method \[\frac{21}{41} \times393000=201293\] \[\frac{20}{41} \times 393000 =191707\]

OpenStudy (waheguru):

wow thanks